# When Large Language Models Meet Vector Databases: A Survey

### v1

@@ Abstract
The recent burst in Large Language Models has opened new frontiers in human-like text processing and generation. However, alongside their remarkable growth, Large Language Models have encountered critical challenges including issues of hallucination, bias, real-time knowledge updates, and the high costs of implementation and maintenance in commercial settings. Vector Databases, another increasingly popular tool, offer potential solutions to these challenges. These databases are adept at handling high-dimensional data and are crucial for tasks such as efficient information retrieval and semantic search. By integrating with Large Language Models, they significantly enhance AI systems’ ability to manage and utilize diverse data more effectively. This survey paper provides an in-depth and unique analysis of the intersection between Large Language Models and Vector Databases.
@@ 1 Introduction
The concept of Artificial Intelligence (AI) was envisioned decades ago, but it is not until ChatGPT unlocked the power of interactive smart AI assistants that people are offered with the unprecedentedly tangible experience of talking to a machine program that ”seems” more knowledgeable than an ordinary person. The phenomenal heat of ChatGPT inherits from the success of Large Language Models (LLMs), one major element of AI. LLMs, exemplified by systems like GPT Brown and et al. (2020); Achiam and et al. (2023), BERT Devlin et al. (2018), and Llama Touvron and et al. (2023), can process, understand, and generate human-like text. Thanks to the power of pre-training which involves training a language model on a massive corpus of text data, LLMs can capture the complexities of human languages, including context, idioms, and even cultural references. Their applications range from simple text completion to complex tasks like translation, understanding, and generation, making them invaluable assets in both commercial and research domains.
@@
However, despite the impressive capabilities of LLMs, they are still under the shadows of doubt in many aspects. One major shortcoming is the problem of hallucination, where LLMs generate plausible but factually incorrect or faithfully nonsensical information Huang et al. (2023). Causes behind this problem include 1. Lack of domain knowledge. The fact that LLMs are primarily trained on public datasets Penedo et al. (2023) inevitably leads to a limited ability to answer domain-specific questions that are out of the scope of their internal knowledge. 2. Real-time knowledge updates. Even if the questions are within the learning corpus of LLMs, their answers may still exhibit limitations because the internal knowledge may be outdated when the outside world is dynamic and keeps changing Onoe et al. (2022). 3. Biases. The datasets used to train LLMs are large which may introduce systematic errors Bender et al. (2021). And essentially every dataset can be questioned with biases issues, including imitative falsehoods Bender et al. (2021), duplicating biases Lee et al. (2022), and social biases Venkit et al. (2023). Moreover, a disadvantage of incorporating LLMs commercially is the expensive cost of maintenance. For an average business entity, applying LLMs for business use is barely feasible. It is almost impossible for a non-tech company to customize and train a GPT model of its own because they don’t have the resources and talents to conduct such big of a project Musser (2023), while frequent API calls to third-party LLM providers like OpenAI can be extremely expensive, not to mention there is a very limited number of such providers in certain areas. Additionally, the oblivion problem of LLMs has been of controversial because LLMs are proven to tend to forget.
Structure Encoder Only Encoder-Decoder Decoder Only
LLMs BERT Devlin et al. (2018), RoBERTa Liu et al. (2019), XLM Lample and Conneau (2019), ALBERT Lan and et al. (2019), ELECTRA Clark and et al. (2020), DeBERTa He and et al. (2020) T5 Raffel et al. (2020), BART Lewis and et al. (2019), mT5 Xue and et al. (2020), M2M-100 Fan and et al. (2021), BigBird Zaheer and et al. (2020), ChatGLM Zeng and et al. (2022) GPT-2 Radford and et al. (2019), GPT-3 Brown and et al. (2020), OPT Zhang and et al. (2022), PaLM Chowdhery and et al. (2023), BLOOM Workshop and et al. (2022), MT-NLG Smith and et al. (2022), GLaM Du and et al. (2022), Gopher Rae and et al. (2021), chinchilla Hoffmann and et al. (2022), LaMDA Thoppilan and et al. (2022), LLaMA Touvron and et al. (2023), GPT-4 Achiam and et al. (2023), BloombergGPT Wu and et al. (2023)
Table 1: Summary of LLMs
Another increasingly popular sub-field of AI is AI databases that support vector data storage and its efficient retrieval at scale, also called vector databases (VecDBs). Whatever kind of multi-modal data that LLMs deal with, searching over many vectors is time-consuming, it is recommended by OpenAI and other LLM vendors to use VecDBs, which always provide LLMs and their applications with cost-effective retrievals and scalable data management.
The intersection of Large Language Models and Vector Databases has been and will still be a burgeoning area of study and application, offering exciting possibilities. This synergy allows for the creation of more powerful, efficient, and versatile AI systems, capable of handling a broad spectrum of tasks with enhanced accuracy and speed. It also paves the way for innovative applications in fields like healthcare, finance, education, and entertainment, where AI can deliver tailored solutions and insights.
@@
In the light of lacking papers that introduce LLMs in the view of Vector Databases, this survey aims to picture how Vector Databases can be potential solutions to refine Large Language Models’ known shortcomings in previous works and hope to offer a unique perspective of future directions in the intersection that is fertile of research potentials. This paper develops as below:
  • Section 2 is presented as introduction to the background knowledge of the two major characters of this paper, LLMs and Vector Databases. Section 2.1 offers a brief view of the popular LLMs and accents on their underlying shortcomings, while Section 3.1 depicts the picture of Vector Databases and their unique capabilities in integration with LLMs.
  • In Section 3, we provide a comprehensive summary of how previous research has effectively combined LLMs with Vector Databases. Section 3.2 delves into a detailed blueprint of the Retrieval-Augmented Generation paradigm in which a unique role that Vector Databases play. Section 3.3 and 3.4 showcases other memory-wise benefits that vector databases can bring for LLMs.
  • Section 4 is dedicated to our analysis of the limitations and unexplored areas at the intersection of LLMs and Vector Databases and highlights the potential future research opportunities in this domain.
@@ 2.1 Large Language Models
Over the last half-decade, we have witnessed the groundbreaking success of Large Language Models (LLMs), marking a significant milestone in the field of Natural Language Processing (NLP). LLMs have revolutionized our views on the approaches to understanding and generating human languages by machines. There are two key factors in the evolution of Large Language Models, the development of Neural Network architectures and the superpower of pre-training models on extensive data.
@@ 2.1.1 Development of Language Models
The next pivotal milestone for NLP is the widespread adoption of the Transformer architecture Vaswani et al. (2017) which has set a new standard for language models. The multi-head self-attention modules and cross-attention modules in the encoders and decoders enable the model to capture long-range dependencies, parallel processing, and contextual understanding. Furthermore, the model starts to rapidly grow in size in the Transformer era. More importantly, the Transformer represents a paradigm shift in NLP field, and the success of the Transformer architecture becomes the significant backbone of LLMs, upon which the encoder-only, encoder-decoder, and decoder-only models are built.
Encoder-Only Models
These models are designed to analyze and understand input text. They process the input to create representations, in the form of embedding vectors, that capture the nuances and context of the language. Examples of the models in this category include BERT Devlin et al. (2018) and its derivatives, such as RoBERTa Liu et al. (2019). They excel in tasks that require a deep understanding of language context, such as sentiment analysis, named entity recognition, and question answering where the answers are contained within the given text.
Encoder-Decoder Models
This architecture consists of two parts: an encoder that processes the input text and a decoder that generates the output. These models are particularly effective in tasks that involve transforming an input into a different output, such as machine translation or text summarizing. The Transformer model itself, as we have discussed before, is an example of this category, which has brought astonishing breakthroughs in machine translation. Another significant model is T5 Raffel et al. (2020), which frames all language tasks as a text-to-text problem, showcasing the versatility of the encoder-decoder framework.
Decoder-Only Models
Decoder-only architectures are optimized for generating text. These models take an input and then continue to generate language based on that input. GPT-3 Brown and et al. (2020) and its successor, GPT-4 Achiam and et al. (2023), are quintessential examples of this category. They are particularly adept at tasks that require creative and coherent text generation, such as content creation, storytelling, and even code generation.
We show the LLMs that have received notable attention during the development of language models in Table 1.
3 LLM+VectorDB
Figure 1: Sample: A sample RAG framework that uses vector databases.
@@ 3.0.1 Power of Pre-training
The scaling laws proposed by OpenAI Kaplan and et al. (2020) highlight a critical trend: the scaling of Pretrained Language Models (PLMs), in terms of both model size and data volume, leads to significant improvement on downstream tasks. This is evidenced by the development of increasingly larger PLMs, such as the 1.76-trillion-parameter GPT-4 Achiam and et al. (2023) and the 540-billion-parameter PaLM Chowdhery and et al. (2023). Unlike their smaller predecessors, like the 330-million-parameter BERT Devlin et al. (2018) and 1.5-billion-parameter GPT-2 Radford and et al. (2019), these large-scale models demonstrate unique behaviors and emergent abilities Wei and et al. (2022) in various tasks like zero-shot learning, which is extremely challenging to the small-scale models. Researchers started to call the large-scale models in the NLP field “Large Language Models” or “LLMs” to denote these significantly scaled PLMs, in order to distinguish the outstanding performance and gain more traction in academia.
@@
The superpower of pre-training lies in its ability to provide models with a general understanding of the languages, which is then tailored through additional training, known as fine-tuning, for specific tasks unlike traditional deep learning methods Dong et al. (2021); Liu and et al. (2024). The advantage is that the model doesn’t start from scratch when learning a new task; it builds upon a rich, pre-existing foundation of language understanding. And this enables what is known as transfer learning Yosinski and et al. (2014). Knowledge gained during this initial phase can be transferred to a wide range of tasks, even those that the model was not explicitly trained on.
3.0.2 Are we there yet? (Challenge faced by pure LLMs)
However, the success of LLMs does not come without challenges. Issues such as data biases Raffel et al. (2020), which can lead to skewed or unfair outcomes Li and et al. (2023), are of significant concern. Ethical considerations, including privacy problems Yao et al. (2024) and the potential for misuse of generative content Ganguli et al. (2022). Training language models to follow instructions with human feedback, like toxic content, also receives quite a bit of debate. Researchers are arguing that the power of LLMs is very superficial, and it is the large pre-training data Schaeffer (2023) that helps the models gain excellent performance on most of the benchmarks. Hallucinations Ji and et al. (2023) are also found to be a major problem with LLMs, casting doubt on the reliability of their outputs. Additionally, the computational resources required for training and running these models are substantial, posing environmental and accessibility questions Strubell and et al. (2019). As the sizes of both LLMs and the training datasets keep getting bigger, the cost of training and inferring is heavily influenced. It is estimated that training the 11-billion parameter version of T5 Raffel et al. (2020) costs over 1.3 million dollars for a single run, whereas one round of training GPT-3 Brown and et al. (2020) of 175 billion parameters using a Tesla V100 cloud instance requires costs 4.6 million dollars. On top of that, training a BERT-based language model Devlin et al. (2018) using 8 V100 GPUs for 36 hours and used a total of 37.3 kWh which is estimated to be more energy-consuming than a gallon of gasoline in terms of $CO_{2}$ emission Dodge and et al. (2022).
@@ 3.1 Vector Databases, the V-factor
While LLMs like ChatGPT are relatively new concepts, database management systems (DBMS) have been thoroughly developed and applied in many aspects in the last 60 years of history, well-recognized for their consistent stability and universality for structured data with fixed formats that excel well with computer storage. However, the development and wide application of deep learning models such as convolutional neural networks He and et al. (2016) and transformers Devlin et al. (2018); Su and et al. (2022), enables the embedding of unstructured multi-modal data like images and text mapped into corresponding fixed-length vector representations, which include the high-dimensional semantic features of original data and the semantic similarities are natively represented by distances between vectors, requiring a new type of DBMS that is specifically designed for handling vector data operations, especially vector search and storage.
@@
There are many kinds of databases, as shown in Table 2, whereas vector databases are the only category of databases that natively support diverse unstructured data with efficient storage, indexing, and retrieval. All these operations could be done based on the vector indexes, which are optimally designed collections of vectors deployed together for ANN search. As a result of various blooming applications on the cloud, various data sources are in many different formats and diverse places. Unlike traditional databases that require structured or semi-structured data that must convey a few restrictions and formats, vector databases are purpose-designed for storing the deep learning embedding of various unstructured data that emerged in real-world applications. On the other hand, distinct from traditional DBMS that searches for exact values within databases, vector databases heavily rely on approximate nearest neighbor (ANN) search for vectors, which searches approximate top-k nearest distance neighbors within the high-dimensional base vector data space that do not necessarily require exact matches.
@@
With the unstructured base data embedded into vectors with high dimensionality, calculating k-nearest neighbors of a given query vector data can be expensive, since it requires distance computation to every point in the dataset and maintaining the top-k results. Such operation would direct to a time complexity of $O(dN+N\log{k})$, where $d$ is the dimensionality and $N$ is the number of vectors, for searching top-k results using pair-wise distance calculation and a heap to keep top-k results.
@@
This calls for a more efficient search technique with satisfying accuracy, the ANN Benchmarks  [^1] https://ann-benchmarks.com/ has showcased the great performance gap between brute-force ANN search and index-enhanced ANN search. Since brute force search is both time-consuming and computationally expensive, vector indexing can solve both of these problems, which optimized for ANN search within VDBMS include tree-based Muja and Lowe (2009); Tao et al. (2009), hash-based Andoni and Indyk (2008), Product Quantization (PQ) Jégou et al. (2011), and graph-based methods Malkov and Yashunin (2018). Among all these indexes, the graph-based Hierarchical Navigable Small Worlds (HNSW) provides state-of-the-art performance and capability with great universality and is widely used within most vector database management systems, including the examples we mentioned in Table 3.
Category Supported Data type Examples
Relational Databases (RDBMS) Structured MySQL, PostgreSQL, Oracle
Document Databases Semi-Structured MongoDB, Couchbase
Time-Series Databases Structured InfluxDB, TimescaleDB
Graph Databases Structured Neo4j, Amazon Neptune
Vector Databases (VDBMS) Unstructured Pinecone, Milvus, Jina, Qdrant
Table 2: Different kinds of mainstream business-level databases.
Name Supported Data Supported Query Vector Index
(Version with year) Vec. Dim. Type Filter Multi-Vector Graph IVF
ChormaDB (2022) 1536 Vec. $\checkmark$ $\checkmark$ $\checkmark$
Manu (2022) 32768 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Milvus (2021) 32768 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Weaviate (2019) 65535 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Qdrant (2021) 4096 Vec. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AnalyticDB-V (2020) $\geq$512 Rel. $\checkmark$ $\checkmark$ $\checkmark$
MongoDB Atlas (v6.0, 2023) 2048 NoSQL $\checkmark$ $\checkmark$
MyScale (2023) 1536 Rel. $\checkmark$ $\checkmark$
Table 3: Comparison of mainstream all-level databases supporting vector data (the version number represents the first version that supports vector search and its release date)

### v2

@@ Abstract
This survey explores the synergistic potential of Large Language Models (LLMs) and Vector Databases (VecDBs), a burgeoning but rapidly evolving research area. With the proliferation of LLMs comes a host of challenges, including hallucinations, outdated knowledge, prohibitive commercial application costs, and memory issues. VecDBs emerge as a compelling solution to these issues by offering an efficient means to store, retrieve, and manage the high-dimensional vector representations intrinsic to LLM operations. Through this nuanced review, we delineate the foundational principles of LLMs and VecDBs and critically analyze their integration’s impact on enhancing LLM functionalities. This discourse extends into a discussion on the speculative future developments in this domain, aiming to catalyze further research into optimizing the confluence of LLMs and VecDBs for advanced data handling and knowledge extraction capabilities.
@@ 1 Introduction
The field of Natural Language Processing (NLP) achieves a great milestone when ChatGPT unlocks the power of interactive smart assistants and people are offered the unprecedentedly tangible experience of talking to a machine program that is seemingly more knowledgeable than an ordinary person. This remarkable breakthrough is largely attributed to the advancements in LLMs, such as GPT  Brown et al. (2020); Achiam et al. (2023), T5  Raffel et al. (2020), and Llama  Touvron et al. (2023). They can process, understand, and generate human-like text. Thanks to the power of pre-training, which involves training a language model on a massive corpus of text data, LLMs can capture the complexities of human languages, including context, idioms, and even cultural references.
@@
However, despite the impressive capabilities of LLMs, they are still under the shadows of doubt in certain aspects. 1. One major shortcoming is the problem of hallucination, where LLMs generate plausible but factually incorrect or faithfully nonsensical information  Huang et al. (2023). There are many potential causes behind this. First, LLMs lack domain knowledge. The fact that LLMs are primarily trained on public datasets  Penedo et al. (2023) inevitably leads to a limited ability to answer domain-specific questions that are out of the scope of their internal knowledge. Moreover, real-time knowledge updates are challenging for LLMs. Even if the questions are within the learning corpus of LLMs, their answers may still exhibit limitations because the internal knowledge may be outdated when the outside world is dynamic and keeps changing  Onoe et al. (2022). Last but not least, LLMs are found to be biased. The datasets used to train LLMs are large, which may introduce systematic errors  Bender et al. (2021). Essentially, every dataset can be questioned with bias issues, including imitative falsehoods  Lin et al. (2022), duplicating biases  Lee et al. (2022), and social biases  Narayanan Venkit et al. (2023). 2. Another disadvantage of incorporating LLMs commercially is the expensive cost Schwartz et al. (2020). For an average business entity, applying LLMs for business use is barely feasible. It is almost impossible for a non-tech company to customize and train a GPT model of its own because they don’t have the resources and talents to conduct such a big project  Musser (2023), while frequent API calls to third-party LLM providers like OpenAI can be extremely expensive, not to mention there are a very limited number of such providers in certain areas. 3. The oblivion problem of LLMs has been controversial because LLMs are found to tend to forget information from previous inputs. It is studied that LLMs also exhibit the behavior of catastrophic forgetting  Luo et al. (2023) as neural networks do  Kemker et al. (2017).
While generative models like LLMs use vector data embeddings to represent unstructured data in our real world, users need a reliable data system to manage and retrieve a considerable amount of data with cost-efficiency. Yet another increasingly popular field that seems orthogonal to LLMs is purpose-built databases for AI that support vector data storage and its efficient retrieval at scale, also known as VecDB. With efficient combinations of VecDBs and LLMs, VecDBs can either be incorporated as an external knowledge base with efficient retrieval providing domain-specific knowledge for LLMs, a memory for LLMs saving previous related chat contents for each user’s dialog tab, or a semantic cache, the aforementioned problems of LLMs can be seamlessly solved.
@@
In the light of lacking papers that introduce LLMs in the view of VecDB, this survey aims to picture how VecDBs can be potential solutions to refine LLMs’ known shortcomings in previous works and hopes to offer a unique perspective of future directions in the intersection that is fertile of research opportunities.
Name Supported Data Supported Query Vector Index
(Version with year) Vec. Dim. Type Filter Multi-Vec. Graph IVF
ChromaDB (2022) 1536 Vec. $\checkmark$ $\checkmark$
Manu (2022) 32768 Vec.+ Ftx. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Milvus (2021) 32768 Vec.+ Ftx. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Weaviate (2019) 65535 Vec.+ Ftx. $\checkmark$ $\checkmark$ $\checkmark$
Qdrant (2021) 4096 Vec.+ Ftx. $\checkmark$ $\checkmark$ $\checkmark$
AnalyticDB-V (2020) $\geq$512 Rel.+Ftx. $\checkmark$ $\checkmark$ $\checkmark$
MongoDB Atlas (v6.0, 2023) 2048 NoSQL+ Ftx. $\checkmark$ $\checkmark$
MyScale (2023) 1536 Rel. $\checkmark$ $\checkmark$
Table 1: Comparison of mainstream all-level databases supporting vector data (the version number represents the first version that supports vector search and its release date). Abbreviations note: Ftx., Rel., Vec. stand for full-text search, relational database, and vector search.
@@ 2.1 Large Language Models
Over the last half-decade, we have witnessed the groundbreaking success of LLMs, marking a significant milestone in the field of NLP. LLMs have revolutionized our views on the approaches to understanding and generating human languages by machines.
@@ 2.1.1 Development of Language Models
The next pivotal milestone for language models is the widespread adoption of the Transformer architecture  Vaswani et al. (2017) which has set a new standard for language models. The multi-head self-attention modules and cross-attention modules in the encoders and decoders enable the model to capture long-range dependencies, parallel processing, and contextual understanding. Furthermore, the model starts to rapidly grow in size in the Transformer era. More importantly, the Transformer represents a paradigm shift in the NLP field, and the success of the Transformer architecture becomes the significant backbone of LLMs.
@@ 2.1.2 Power of Pre-training
The scaling laws proposed by OpenAI  Kaplan et al. (2020) highlight a critical trend: the scaling of Pretrained Language Models (PLMs), in terms of both model size and data volume, leads to significant improvement in downstream tasks. This is evidenced by the development of increasingly larger PLMs, such as the 1.76-trillion-parameter GPT-4  Achiam et al. (2023) and the 540-billion-parameter PaLM  Chowdhery et al. (2023). Unlike their smaller predecessors, like the 330-million-parameter BERT  Devlin et al. (2018) and 1.5-billion-parameter GPT-2  Radford et al. (2019), these large-scale models demonstrate unique behaviors and emergent abilities  Wei et al. (2022) in various tasks like zero-shot learning, which is extremely challenging for the small-scale models.
@@
The power of pre-training lies in its ability to provide models with a general understanding of the languages, which is then tailored through additional training, known as fine-tuning, for specific tasks. The advantage is that the model doesn’t start from scratch when learning a new task; it builds upon a rich, pre-existing foundation of language understanding. This enables what is known as transfer learning  Yosinski et al. (2014).
2.1.3 Are We There Yet? Challenges Faced by Pure LLMs
Although LLMs have achieved remarkable success, they face significant challenges that cannot be overlooked. Three primary challenges are particularly noteworthy in the context of pure LLMs. Addressing these challenges is crucial for the continued advancement and practical application of these models. We have outlined these challenges as follows:
  • Model Editing involves integrating more flexible architectures into models, enabling targeted updates or edits without the need for complete retraining. This presents a complex challenge: alterations in one part of the model can unexpectedly affect other areas. Ensuring that these updates neither degrade performance nor introduce biases is a difficult task. Model editing also includes elements such as personality, emotions, opinions, and beliefs. While these aspects have received some attention, they largely remain unexplored territories  Yao et al. (2023). Ethical concerns, including data biases  Raffel et al. (2020), privacy issues  Li et al. (2023a), and the potential misuse of generative content  Ganguli et al. (2022), are significant challenges in model editing.
  • Hallucinations in LLMs  Ji et al. (2023) pose a major challenge, undermining the reliability of their outputs. LLMs often generate plausible yet factually incorrect or nonsensical information  Huang et al. (2023). Several factors contribute to this issue. First, LLMs lack domain-specific knowledge, primarily due to their training on public datasets  Penedo et al. (2023), limiting their ability to answer domain-specific questions. Additionally, updating LLMs with real-time knowledge is a challenge, leading to outdated internal knowledge in a rapidly changing world  Onoe et al. (2022).
  • Substantial computational resources required for training and operating LLMs raise environmental and accessibility concerns  Strubell et al. (2019), contradicting the principles of sustainable AI  Wu and Raghavendra (2022). The escalating sizes of LLMs and their training datasets significantly influence training and inference costs. The environmental footprint of these technologies is substantial, with significant greenhouse gas emissions and energy consumption during model training and tuning  Kaack et al. (2022). This situation necessitates a critical evaluation of AI’s environmental impact, encouraging the development of more energy-efficient models and the adoption of green computing practices  Wynsberghe (2021).
Transitioning to VecDBs could potentially comprehensively address these challenges. VecDBs, with their efficient data structuring and retrieval capabilities, offer a promising solution for enhancing model editing flexibility, reducing hallucinations by improving domain-specific responses, and optimizing computational resource utilization, thereby aligning more closely with the principles of sustainable AI.
@@ 2.2.1 A distinct database optimized for vector data
While LLMs like ChatGPT are relatively new concepts, database management systems (DBMS) have been thoroughly developed and applied in many aspects in the last 60 years of history and are well-recognized for their consistent stability and universality for structured data with fixed formats that excel well with computer storage. However, the development and wide application of deep learning models such as convolutional neural networks He et al. (2016) and transformers Devlin et al. (2018), enable the embedding of unstructured multi-modal data like images and text mapped into corresponding fixed-length vector representations, which include the high-dimensional semantic features of original data, and the semantic similarities are naturally represented by distances between vectors, requiring a new type of DBMS that is specifically designed for handling vector data operations, especially search and storage.
@@
There are many kinds of databases, whereas VecDBs are the only category of databases that naturally support diverse unstructured data with efficient storage, indexing, and retrieval. As a result of various blooming applications on the cloud, various data sources exist in many different formats and diverse places. Unlike traditional databases that require structured or semi-structured data that must convey a few restrictions and formats, VecDBs are purpose-designed for storing the deep learning embedding of various unstructured data that has emerged in real-world applications. On the other hand, distinct from traditional DBMSs that search for exact values within databases, the semantic feature of vectors in VecDBs needs an approximate search for vectors, which searches approximate top-k nearest-distance neighbors within the high-dimensional base vector data space that do not necessarily require exact matches, representing top-k semantically close data.
2.2.2 Efficient vector retrieval within VDBMS
@@
With the unstructured base data embedded into vectors with high dimensionality, calculating the k-nearest neighbors of a given query vector data can be expensive since it requires distance computation to every point in the dataset and maintaining the top-k results. Such an operation would result in a time complexity of $O(dN+N\log{k})$, where $d$ is the dimensionality and NN is the number of vectors, for exhaustively searching top-k results using pair-wise distance calculation and a heap to keep top-k results.
@@
Since brute-force search is time-consuming and computationally expensive, this calls for a more efficient search and storage technique with satisfying accuracy. Vector indexing can solve both of these problems, which optimized for ANN search within VDBMS include tree-based Muja and Lowe (2009); Tao et al. (2009), hash-based Andoni and Indyk (2008), Product Quantization (PQ) Jégou et al. (2011), and graph-based methods Malkov and Yashunin (2018), while the ANN Benchmarks  [^1] https://ann-benchmarks.com/ has also showcased the great performance gap between brute-force search and index-enhanced ANN search. All these operations could be efficiently done based on the vector indexes, which are optimally designed collections of vectors deployed together for approximate nearest neighbor (ANN) search. The ANN Benchmarks have also showcased the great performance gap between brute-force search and index-enhanced ANN search. Also, according to them, among all these indexes, the graph-based Hierarchical Navigable Small Worlds (HNSW) provides state-of-the-art performance and capability with great universality and is widely used within most VecDB management systems, including the examples we mentioned in Table 1.
2.2.3 A feasible solution of knowledge base for LLMs
Traditional full-text and keyword search engines, such as Elastic Search and Amazon OpenSearch, are based on term frequency metrics like BM25, which has proved practical for many search applications. However, such search techniques require significant investment in time and expertise to tune them to account for the meaning or relevance of the terms searched. On the other hand, they lack multi-modal storage and retrieval support, as VecDBs do.